[Hallc_running] [New Logentry] RC daily update

simona at jlab.org simona at jlab.org
Sun Oct 7 15:15:01 EDT 2018

Logentry Text:
Yesterday after the recovery of the ion pump at 2L23 we got beam back around 11 am and we resumed production on Al dummy at  Q2=4.4 W=2.74 x=0.4 theta_Shms=9.81. This was interrupted by a beam transport issue that arose when recovering from an RF trip [https://logbooks.jlab.org/entry/3605153] , [https://logbooks.jlab.org/entry/3605144], [https://logbooks.jlab.org/entry/3605184]. Once the magnet that caused the issue was put back at the proper current beam was resumed to the halls. We managed to finish the Al running at about 1 pm and changed SHMS angle to take data at Q2=4.4 W=2.74 x=0.4 theta_Shms=12.81 on Al.

At about this time all RF separators tripped and the saga of the RF separators recovery started [https://logbooks.jlab.org/entry/3605189]. Pass 5 separator was up sooner and we managed to take Al dummy data at the new SHMS angle and also some LH2 data starting with 5 pm. The Ops experts were on site trying to recover the pass 2 separator. Initially it was believed that a tube in IOT 1 that operates Pass 2 separator was bad so a reconfiguration was done to use IOT 2 in place of IOT 1 to operate Pass 2. Beam was sent to all halls for a little while but there were many BLM and IC trips. At about 5 am the decision was made to turn off the pass 2 separator (no beam to Hall A) and focus on sending beam to B and C but then the 5th pass separator tripped and experts were called on site again [https://logbooks.jlab.org/entry/3605383]. At about the same time the familiar PSS glitch put the NL in power permit. 

In conclusion we got very little physics beam since yesterday and we are pretty far behind with the data taking. We are still at Q2=4.4 W=2.74 x=0.4 theta_Shms=12.81 with 33 % completed of the marginal 58% we were shooting for.

As of now the Ops experts found the problem with the RF separator: it resided with the IOT power supply [https://logbooks.jlab.org/entry/3605488]. A burned lead and stand-offs were found due to a loose connection that overheated and arced. This was fixed during the day shift.  Also the NL03 zone that dropped earlier this morning was recovered. As Arne put it, all systems are GO [https://logbooks.jlab.org/entry/3605492].

As of now we are waiting for beam. When it comes we will finish Q2=4.4 W=2.74 x=0.4 theta_Shms=12.81 LH2 running and then move to the last SHMS angle at this kinematics.

For a summary of the data taken so far check out this link [https://hallcweb.jlab.org/wiki/index.php/Kaon_LT_Status_Page].

There will be no RC meeting today.

P.S. New issue: [https://logbooks.jlab.org/entry/3605541]


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3605542
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181007/95672c50/attachment.html>

More information about the Hallc_running mailing list